OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form

OpenStudy (ivettef365):

y = mx + b

OpenStudy (ivettef365):

Therefore y - 4 = -2x + 10 Y = -2x + 14

OpenStudy (anonymous):

Thanks :)